AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 00:42:53 2025
Author
*
:
Subject
*
:
<p><strong>20:00</strong></p> <p>The beam is set at the correct energy, operators cross-checked it again. What we see at 750 is due to the molecular beam. The operotors are trying to optimize and get rid of this in the next 30min.</p> <p>Temps and system checks OK</p> <p>Leakage current OK</p> <p>Merger and tape server OK</p> <p>Potentiometer at 28.06 mm and 29.23 mm</p> <p><strong>20:30</strong></p> <p>Keep going with the beam tests. A new peak appears around 950 keV( See attached). Still try to understand what it is. The operators keep going</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown